# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

### v3

@@ Abstract
Long-duration talking video synthesis faces persistent challenges in simultaneously achieving high video quality, portrait and temporal consistency, and computational efficiency. As video length increases, issues such as visual degradation, loss of identity consistency, temporal incoherence, and error accumulation become increasingly prominent, severely impacting the realism and reliability of generated results. To address these issues, we present LetsTalk, a diffusion transformer framework that incorporates multimodal guidance and a novel memory bank mechanism, explicitly maintaining contextual continuity and enabling robust, high-quality, and efficient long-duration talking video generation. Specifically, LetsTalk introduces a memory bank combined with a noise-regularized training strategy to mitigate error accumulation and sampling artifacts during long video generation. To further enhance efficiency and spatiotemporal consistency, LetsTalk employs a deep compression autoencoder and a spatiotemporal-aware transformer with linear attention for effective multimodal fusion. Furthermore, we systematically analyze three multimodal fusion schemes, adopting deep (Symbiotic Fusion) for portrait features to ensure visual consistency, and shallow (Direct Fusion) for audio to synchronize animation with speech while preserving motion diversity. Extensive experiments demonstrate that LetsTalk achieves state-of-the-art generation quality, producing temporally coherent and realistic talking videos with enhanced diversity and liveliness, while maintaining remarkable efficiency with 8 fewer parameters than previous approaches.
Index Terms:
Talking Head Generation, Long-duration Generation, Multimodal Guidance, Linear Diffusion Transformer, Deep Compression Autoencoder
@@ I Introduction
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc. Despite recent progress, long-duration talking video synthesis still faces persistent challenges in simultaneously achieving high video quality, portrait and temporal consistency, and computational efficiency. As video length increases, issues such as visual degradation, loss of identity consistency, temporal incoherence, and error accumulation become increasingly prominent, severely impacting the realism and reliability of generated results.
Previous works [1, 2] use a 3D morphable model (3DMM) [3, 4, 5] to represent reference portraits and drive them with audio signals. These approaches input driven 3DMM parameters (e.g., coefficients, poses) into neural renderers, achieving temporal and portrait consistency yet lacking output diversity, often showing static features like unblinking eyes and rigid hair. In contrast, audio-to-2D facial motion methods [6, 7, 8, 9, 10] directly align audio with facial dynamics via latent space learning. However, their limited disentanglement of static and dynamic components compromises both temporal consistency and animation diversity.
Recently, diffusion models [11, 12, 13] have emerged as a powerful paradigm for generating realistic 2D images [13, 14] and have demonstrated remarkable success in video generation [15, 16]. Notably, AnimateAnyone [17] introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion. Building on this, recent methods [18, 19, 20] have extended this approach by replacing pose signals with audio signals for talking head generation. These methods leverage diffusion models to generate diverse unconscious actions, significantly enhancing the richness of generated content.
However, several critical challenges remain unaddressed, particularly in the context of long-duration video generation. As the length of generated videos increases, existing methods struggle to maintain high visual quality, portrait consistency, and temporal coherence. Progressive accumulation of residual noise and artifacts (e.g., facial distortions or flickering) becomes increasingly severe over extended sequences, ultimately degrading output fidelity and realism. In addition to these long-duration challenges, current approaches also suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific tasks; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity. Furthermore, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio.
@@
In this work, we introduce LetsTalk, a diffusion transformer framework designed for efficient and high-quality generation of realistic, long-duration talking videos under multimodal conditions. A key component for long video generation in LetsTalk is the memory bank mechanism, which enables each generated clip to access information from previous clips, ensuring smooth transitions and contextual continuity. To address the discrepancy between training and inference, we propose a noise-regularized training strategy that explicitly simulates inference-time noise accumulation, thereby mitigating error accumulation and sampling artifacts as the video extends.
To further enhance efficiency and scalability, LetsTalk employs a deep compression autoencoder that reduces input frames by 32 in the latent space, significantly decreasing the number of tokens while preserving essential visual details. Each transformer block is equipped with dedicated temporal and spatial modules, leveraging linear attention [21] to efficiently model both intra-frame and cross-frame coherence without incurring excessive computational cost. This architecture enables LetsTalk to maintain high generation quality and temporal consistency even for long video sequences.
@@
Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference portrait and driving audio. To better exploit multimodality for guiding long-duration video generation, we first summarize three fusion schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion. We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities. We further investigate the underlying factors influencing such adaptability, thereby bridging theoretical gaps in conditional guidance modeling. To effectively guide long-duration video generation, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency throughout extended sequences due to its visual similarity to the video. For audio, we adopt Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion) over long video durations.
We summarize the contributions of this work as follows
@@
We propose LetsTalk, a novel diffusion transformer framework for long-duration talking video generation, which simultaneously achieves high video quality, portrait and temporal consistency, and computational efficiency.
We introduce a memory bank mechanism combined with a noise-regularized training strategy, which explicitly maintains contextual continuity and effectively mitigates error accumulation and sampling artifacts during long video generation.
We design an efficient spatiotemporal modeling architecture, leveraging a deep compression autoencoder and linear attention, which achieves comparable visual fidelity with only 1/8 the parameters of previous methods.
@@
We systematically analyze multimodal fusion schemes and adopt deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio, enabling LetsTalk to achieve state-of-the-art visual quality and temporal coherence in talking video generation.
@@ II-B Audio-driven Portrait Image Animation
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert [7] achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet [6] and MakeItTalk [8] introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker [1], Genface [30], and DiffTalk [31] further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads [32], DreamTalk [33], and VividTalk [34], which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 [35], EMO [36] and AniPortrait [2] have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker [37] introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait [38] extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo [18, 20], EchoMimic [19] and X-Portrait [39] have adopted the latent diffusion framework pioneered in AnimateAnyone [17], adding more control over elements such as lip movements, expressions, and poses.
@@ III Methodology
We present LetsTalk, a diffusion transformer for audio-driven portrait animation, as illustrated in Fig. 2. This section first introduces the diffusion formulation and transformer backbone in Section III-A. We then detail our long-range consistency mechanism, including the memory bank and noise-regularized training, in Section III-C. Next, we describe the efficient spatiotemporal transformer backbone in Section III-B. Finally, we analyze multimodal fusion schemes and our guidance strategy in Section III-D and Section III-E.
III-A Preliminaries
Diffusion Formulation The diffusion process incrementally corrupts data with Gaussian noise: , where are predetermined hyperparameters. By reparameterizing, we obtain: , where , forming a Markov chain over steps.
Diffusion models are trained to learn the reverse process, modeled as: , with neural networks parameterizing . This is optimized via the variational lower bound of :
With both and Gaussian, is computed from their means and covariances. Reparameterizing as a noise-prediction network enables training via mean-squared error between and sampled noise :
To incorporate a learnable reverse covariance , we minimize the full term, following Nichol and Dhariwal: optimizing with and with the full .
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) [49] in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
III-B Efficient SpatioTemporal Transformer with Linear Attention
Deep Compression Autoencoder: We build upon latent diffusion models and employ a deep compression autoencoder [21] that downsamples images by a factor of 32, maintaining reconstruction quality comparable to SDXL’s autoencoder. This design leverages the redundancy in high-resolution video frames and decouples compression from denoising, allowing the autoencoder to preserve crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment. Compared to conventional autoencoders, this approach significantly reduces memory and computation requirements while maintaining high fidelity.
The compressed latent feature is divided into non-overlapping patches, with spatial-temporal positional embeddings [50] to maintain structural relationships. This compact representation enables efficient transformer processing and reduces memory requirements by 4 compared to conventional approaches.
Interleaved Fusion with Linear Attention: Given a sequence of latent features , we employ a transformer architecture to predict the noise for removal. Unlike the original DiT [29], which processes each frame independently, our framework introduces interleaved spatiotemporal fusion to jointly model spatial and temporal dependencies.
A naive approach would flatten into tokens and apply standard self-attention, resulting in prohibitive quadratic complexity for high-resolution videos. To address this, we adopt a dual-block decomposition strategy that alternates between spatial and temporal attention blocks, significantly reducing the computational cost to .
For spatial modeling, we reshape the input to and apply ReLU-based linear attention [51] within each frame:
where and .
For temporal modeling, we transpose the tensor to and apply standard softmax attention across frames, allowing us to preserve the expressive power of full attention for inter-frame interactions.
By interleaving these spatial and temporal attention blocks, our architecture achieves near-linear efficiency while maintaining a global receptive field, enabling scalable and effective modeling of high-resolution, long-duration talking videos.
@@ III-B Efficient SpatioTemporal Transformer with Linear Attention
Mix-FFN block In spatial modeling, linear attention converges slower despite matching final performance. Following SANA [21], we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU. The depth-wise convolution enhances local feature capture, compensating for ReLU linear attention’s weaker locality. This enhances local feature capture and compensates for the weaker locality of linear attention.
Adaptive Integration Following VDT [16], we insert an adaptive fully-connected layer between attention layers to fuse temporal dynamics into spatial representations:
where is a learnable scalar and denotes a fully-connected layer. This enables adaptive fusion of temporal dynamics into spatial representations.
@@
III-C Memory Bank and Noise-Regularized Training for Long-Term Consistency
@@
III-D Analysis of Multimodal Fusion Schemes
@@
Modality Adaptation. The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks. Direct Fusion, though offering weak alignment, is highly cost-effective for modalities with significant differences, such as text and image, making it ideal for resource-constrained scenarios. Siamese Fusion offers enhanced design flexibility by allowing adjustable modality alignment. This is exemplified in models such as ControlNet [14], which employs visual cue-level alignment, and AnimateAnyone [17], which utilizes pixel-level alignment. The strength of alignment can thus be precisely tailored to meet the specific requirements of the task at hand. Symbiotic Fusion enhances parameter efficiency compared to Siamese Fusion, achieving strong alignment with fewer parameters. This makes it particularly effective for large-scale applications, such as AnimateAnyone [17], where shared parameters ensure robust and efficient multimodal integration. The comparative analysis of the three fusion approaches is summarized in Table I.
@@
III-E Guidance on Reference Portrait and Audio
Given the reference image, we require that the character of the generated video is consistent with the reference. Since the image modality is close to the video, Symbiotic Fusion proves appropriate. Specifically, we first encode the reference image and obtain the portrait feature in the latent space. Then we temporally repeat the portrait feature with frames and concatenate it with the latent feature along the spatial dimension, which can be formulated as: , where denotes the temporally extended portrait feature. This operation yields a concatenated latent feature . By leveraging shared self-attention in the fused feature space, this method maintains consistent character representation across frames. In contrast, Direct Fusion uses shallow portrait features, leading to misalignment in deeper layers. While Siamese Fusion mitigates this by maintaining parallel transformer branches, it increases computational complexity without significant advantages over Symbiotic Fusion.
As for the audio, we require it to reasonably drive the animation, especially the lip motion [1]. However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to increase diversity. Considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio.
For the audio representation, to accommodate the tokenization mechanism of the transformer, we follow [18] and extract richer semantic information from different layers of the wav2vec encoder [54]. Through nonlinear mapping, we obtain audio tokens enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Section IV-F verify the efficacy of our mechanism in feature fusion.

### v4

@@ Abstract
Long-duration talking video synthesis faces enduring challenges in achieving high video quality, portrait and temporal consistency, and computational efficiency. As video length increases, issues such as visual degradation, identity inconsistency, temporal incoherence, and error accumulation become increasingly problematic, severely affecting the realism and reliability of the results. To address these challenges, we present LetsTalk, a diffusion transformer framework equipped with multimodal guidance and a novel memory bank mechanism, explicitly maintaining contextual continuity and enabling robust, high-quality, and efficient generation of long-duration talking videos. In particular, LetsTalk introduces a noise-regularized memory bank to alleviate error accumulation and sampling artifacts during extended video generation. To further improve efficiency and spatiotemporal consistency, LetsTalk employs a deep compression autoencoder and a spatiotemporal-aware transformer with linear attention for effective multimodal fusion. We systematically analyze three fusion schemes and show that combining deep (Symbiotic Fusion) for portrait features and shallow (Direct Fusion) for audio achieves superior visual realism and precise speech-driven motion, while preserving diversity of movements. Extensive experiments demonstrate that LetsTalk establishes new state-of-the-art in generation quality, producing temporally coherent and realistic talking videos with enhanced diversity and liveliness, and maintains remarkable efficiency with 8 fewer parameters than previous approaches.
@@ I Introduction
Animating a portrait image with speech, commonly referred to as talking head generation, aims to synthesize a speaking portrait from a single image and accompanying audio. This technology has promising applications in gaming, film production, and digital marketing. Despite recent advancements, long-duration talking video synthesis continues to grapple with maintaining high video quality, portrait consistency, temporal coherence, and computational efficiency [1, 2, 3, 4]. As video duration increases, issues such as visual degradation, loss of identity, and error accumulation become more pronounced, diminishing overall realism and reliability.
Traditional approaches mainly rely on 3D morphable models (3DMM) [5, 6, 7], which deliver temporal and identity consistency but are often limited in their expressiveness [1, 2]. Audio-to-2D facial motion techniques [8, 9, 10, 11, 12], along with multimodal guidance strategies such as Anyonenet [3] and TalkCLIP [4], enrich animation diversity and introduce control via text or voice. Recent progress further enables personalized head movement [13] and text-audio-guided video generation [14], but existing methods often struggle to effectively disentangle static and dynamic factors, thereby sometimes compromising consistency or limiting expressiveness across longer time spans.
Diffusion models [15, 16, 17] have recently achieved remarkable success in generating realistic 2D images [17, 18] and videos [19, 20]. AnimateAnyone [21] showcases the strength of motion modules and ReferenceNet for pose-controllable human animation within diffusion frameworks. Building on these advances, recent works [22, 23, 24] directly employ audio as guidance to synthesize talking head videos, leveraging diffusion models to improve both diversity and expressiveness.
Nevertheless, significant challenges remain for long-duration talking head generation. As video length increases, existing methods frequently struggle to preserve visual quality, portrait consistency, and temporal coherence. Accumulated noise and artifacts, such as distortions and flickering, further undermine realism. Efficiency is often constrained by reliance on auxiliary spatial signals for audio control, unnecessary parameter overhead from fine-tuning generic models for animation tasks, and computationally intensive architectures, including low-compression autoencoders and standard attention schemes. Additionally, effective strategies for fusing multimodal conditions, such as reference portraits and driving audio, have not been sufficiently explored in conditional video generation frameworks.
@@
In this work, we introduce LetsTalk, a diffusion transformer framework crafted for efficient, high-quality generation of realistic, long-duration talking videos under multimodal conditions. A central innovation in LetsTalk for long video generation is the Noise-Regularized Memory Bank, which enables each generated clip to access information from prior clips while explicitly simulating inference-time noise accumulation during training. This mechanism effectively preserves contextual continuity and mitigates error accumulation and sampling artifacts throughout extended video sequences, ensuring smooth transitions and high temporal consistency. To further improve efficiency and scalability, LetsTalk incorporates a Deep Compression Autoencoder that compresses input frames by 32 in the latent space, drastically reducing the number of tokens while retaining essential visual features. Transformer blocks are equipped with specialized temporal and spatial modules, utilizing Linear Attention [25] for efficient modeling of both intra-frame and inter-frame coherence without excessive computational cost. This design empowers LetsTalk to sustain high generation quality and temporal consistency over long video durations.
@@
For multimodal guidance, defined here as fusing reference portraits and driving audio for video generation, we first categorize three fusion schemes by their integration tightness: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion. These schemes are analyzed in terms of integration depth, conditional state, and parameter sharing, providing a systematic summary of their adaptability to diverse modalities. We further examine the underlying factors that influence such adaptability, addressing theoretical gaps in conditional guidance modeling. For effective long-duration video generation, we employ Symbiotic Fusion with the reference image to ensure strong portrait consistency across extended sequences due to its visual proximity to the generated video, and use Direct Fusion for audio to tightly synchronize lip movement and expressions with speech while fostering diversity in other movements (e.g., blinking, head turns) across longer durations.
We summarize the contributions of this work as follows:
@@
We present LetsTalk, a novel and efficient diffusion transformer framework for scalable, long-duration talking video generation, featuring a powerful spatiotemporal modeling architecture that achieves high video quality and computational efficiency with significantly fewer parameters.
We propose a noise-regularized memory bank mechanism that preserves contextual continuity and effectively curbs error accumulation, substantially improving temporal consistency in extended video sequences.
@@
We provide a systematic analysis of multimodal fusion schemes and demonstrate that combining deep Symbiotic Fusion for portrait features and Direct Fusion for audio achieves superior visual quality and temporal coherence.
Experiments demonstrate that our method achieves realistic, coherent, and diverse long-duration talking videos with optimal quality-efficiency trade-offs.
@@ II-B Audio-driven Portrait Image Animation
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization and frame generation [9, 8, 10]. The integration of 3D information and more expressive head movements in models such as SadTalker [1] and Genface [34] further improved realism. A major breakthrough comes from the adoption of diffusion models, exemplified by works like Diffused Heads [35], DreamTalk [36], and VividTalk [37], which produce high-fidelity talking head videos with enhanced lip synchronization and diverse expressions. Recent innovations have explored multimodal guidance and personalized control mechanisms. Notably, Anyonenet [3] pioneered synchronized speech and head generation for arbitrary identities, enabling voice and video synthesis from a single image. TalkCLIP [4] introduced text-guided expressive speaking styles, allowing users to control facial expressions via natural language. Yi et al. [13] proposed personalized head movement prediction, learning individual motion patterns from short video clips for more natural animation. TA2V [14] further explored text-audio guided video generation, leveraging multi-modal inputs for enhanced synthesis. Very recent works such as VASA-1 [38], EMO [39], and Hallo [22, 24] focus on emotionally expressive talking faces and advanced control over lip movements, expressions, and poses, pushing the boundaries of realism and user-controllability. These approaches, often employing latent diffusion frameworks pioneered in AnimateAnyone [21], demonstrate the rapid progress and increasing capabilities of audio-driven portrait image animation.
@@ III Methodology
We propose LetsTalk, a diffusion transformer framework for audio-driven portrait animation, as depicted in Fig. 2. Section III-A introduces the efficient spatiotemporal transformer backbone. Section III-B details our mechanism for long-range consistency, including the memory bank and noise-regularized training. Section III-C analyzes multimodal fusion schemes, and Section III-D elaborates our multimodal guidance strategy.
III-A Efficient SpatioTemporal Transformer with Linear Attention
Deep Compression Autoencoder. We employ a deep compression autoencoder [25] that downsamples images by 32, achieving reconstruction quality comparable to SDXL while greatly reducing memory and computation costs. Separation of compression and denoising enables preservation of essential temporal cues, such as lip-sync dynamics. Compressed latent features are partitioned into non-overlapping patches and assigned spatial-temporal positional embeddings [50] to enforce structural consistency. This approach facilitates efficient transformer processing and further decreases memory usage by about four times compared to conventional methods.
Interleaved Attention Mechanism. Instead of conventional frame-by-frame processing as in DiT [33], we process latent features with interleaved spatial and temporal attention. Full-token attention incurs quadratic complexity , so we alternate spatial linear attention [51] and temporal softmax attention, reducing total cost to . Specifically, for spatial modeling, we reshape and apply ReLU-based linear attention:
where . For temporal modeling, features are transposed to , then subjected to softmax attention across the temporal dimension. This alternating strategy maintains global context and ensures scalability for high-resolution, long-duration inputs.
@@ III-A Efficient SpatioTemporal Transformer with Linear Attention
Mix-FFN Block. Although linear attention achieves similar final performance, its convergence may be slower. Following SANA [25], we replace the standard MLP-FFN with Mix-FFN, which integrates an inverted residual block, a depth-wise convolution, and a gated linear unit (GLU). This convolution enhances localization ability, compensating for the weaker locality observed with ReLU-based linear attention.
@@
III-B Long-duration Generation with Noise-Regularized Memory Bank
@@
III-C Analysis of Multimodal Fusion Schemes
@@
Modality Adaptation. Fusion schemes must adapt to different modalities for effective multimodal tasks. Direct Fusion offers weak alignment but is highly efficient, making it suitable for modalities with large differences, such as text and image, especially in resource-limited settings. Siamese Fusion provides greater flexibility by allowing tunable alignment strength, as seen in ControlNet [18] (cue-level) and AnimateAnyone [21] (pixel-level), enabling precise control based on task demands. Symbiotic Fusion further improves parameter efficiency over Siamese Fusion, achieving strong alignment with fewer parameters—beneficial for large-scale applications like AnimateAnyone [21]. A comparison of these fusion approaches is shown in Table I.
@@
III-D Guidance on Reference Portrait and Audio
Given the reference image, we require that the character of the generated video is consistent with the reference. Since we adopt Symbiotic Fusion, leveraging the similarity between image and video modalities. The reference image is encoded into a portrait feature , temporally repeated across frames, and concatenated with video latent features to form , where . Shared self-attention in this fused space ensures consistent character representation across frames. In comparison, Direct Fusion leads to misalignment in deep layers, while Siamese Fusion adds complexity without significant benefits over Symbiotic Fusion.
For audio-driven animation, especially for lip motion [1], the modality gap between audio and video motivates the use of Direct Fusion, which is simple and preserves output diversity (e.g., blinking, head movements). Following [22], we extract multi-layer wav2vec [54] features and apply nonlinear mapping to obtain audio tokens enriched with semantic information. These are temporally synchronized with video features to guide animation. Ablation studies in Section IV-F validate the effectiveness of this fusion mechanism.